if a man can earn 250 in one month in 21 months he will earn
 

Grade:7

9 Answers

Smitha Vishwanath
7 years ago
Multiply 21 months ti each months earning(250) 5250.
Thats it.
pratik panchal
7 years ago
They earn 250 per month then in 21 months they earn = 250 × 21= Rs. 5250 ✔ans.It`s so easy if that type of question come in ur exam then multiply given months × rupes of per months
BADUGU VIJAY KUMAR YADAV
7 years ago
it is easy.for 1 month he earns 250,for 2 months how much he earns 2(250)=500.
similarly for 21 months he earn=21(250)=5250 that’s it.i hope you are understand.
badugu vijaya kumar yadav
7 years ago
months*earning permonth=21*250=5250
i hope you are understand.it is very easy.practise some problems related to this problem,you can easily solve it.
Ramana
7 years ago
it is very easy to answer,one month he earn=250
so 21months he will earn=250*21=5250
you are easy to solve this to practice more problems related to this question
utpal raj
7 years ago
in one month he can earn 250 
upto 21 month he has to earn 250+250+250........21 times 
so 250*21
which equals to 5250
REDDYPAVAN
7 years ago
@pallaba
The man earn 250 in one month.for 21 months he earn 21*250=5250rupees.if they give time in years you convert years into months and calculate.
jyoti bhatia
7 years ago
If a man earns 250 rupees in one month, then in 21 months he will earn 21*250 = 5250 rupees. So, the man will earn 5250 rupees in 21 months.
amrita amrita
5 years ago
A man earn 250 in a month.so
1 =250
then in 21 month he will earn 250 multiply 21= 5250.
ANSWER :-5250